a. The points 𝐴,𝐡 and 𝐢 have co-ordinates (βˆ’1,2),(1,1) and (2,3) respectively. Sketch the triangle 𝐴𝐡𝐢 .By calculating the lengths of the sides of this triangle, determine if it is scalene, isosceles, or equilateral.

b. Find the distance 𝑀𝐡, where 𝑀 is the midpoint of 𝐴𝐢, and hence find the area of the triangle 𝐴𝐡𝐢.


1
Expert's answer
2022-02-08T12:28:40-0500


a.


"AB=\sqrt{(1-(-1))^2+(1-2)^2}=\sqrt{5}"

"BC=\sqrt{(2-1)^2+(3-1)^2}=\sqrt{5}"

"AC=\sqrt{(2-(-1))^2+(3-2)^2}=\sqrt{10}"

The triangle ABC is isosceles.


b.


"M(\dfrac{x_A+x_C}{2}, \dfrac{y_A+y_C}{2})"

"M(\dfrac{-1+2}{2}, \dfrac{2+3}{2})"

"M(\dfrac{1}{2}, \dfrac{5}{2})"

"BM=\sqrt{(x_M-x_B)^2+(y_M-y_B)^2}"

"=\sqrt{(\dfrac{1}{2}-1)^2+(\dfrac{5}{2}-1)^2}=\dfrac{\sqrt{10}}{2}(units)"

"AC=\sqrt{(x_C-x_A)^2+(y_C-y_A)^2}"

"=\sqrt{(2+1)^2+(3-2)^2}=\sqrt{10}(units)"

"S_{ABC}=\dfrac{1}{2}BM\cdot AC=\dfrac{1}{2}(\dfrac{\sqrt{10}}{2})(\sqrt{10})"

"=\dfrac{5}{2}({units}^2)"